OpenStudy (anonymous):

URGENT Guy started with the day with some money in his pocket. Then when he took some money out of his bank account, he had $2 more than twice the amount he started with. He then received a birthday present which he doubled the amount of money he had. Guy ended the day with $88. How much money did he start with? Please show all work.

OpenStudy (anonymous):

Hi wangkan07 Can I help you with this one...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oke ;-) Lets cal the moneey he has in the morning = X

OpenStudy (anonymous):

let's cal the amount he gets from the bank = Y

OpenStudy (anonymous):

He then has X + Y in his pocket

OpenStudy (anonymous):

What do they tell you about how much that is?

OpenStudy (anonymous):

the total mount in the bank?

OpenStudy (anonymous):

no... the X + Y that he then has

OpenStudy (anonymous):

they tell you is 2 dollar more then twice the amount he started with

OpenStudy (anonymous):

So X + Y = .... can you fill that in?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

could you give me a hint?

OpenStudy (anonymous):

The amount he starts with is X When he adds the amound from the baqnk (which is Y) he has X + Y They say that that is $2 more than twice the amount he started with twice the amount he started with is 2X $2 more than that is 2X + 2 so X+ Y = 2X + 2 Yes?

OpenStudy (anonymous):

\[\left[\begin{array}{l}X+Y=2X+2 \\2(X+Y)=88 \end{array}\right]\] \[\left[\begin{array}{l}Y=X+2 \\X+Y=44 \end{array}\right]\] \[X+X+2=44\] \[2X=42\] \[X=21\] \[Y=X+2\] \[Y=21+2\] \[Y=23\]